# Bourbaki degree of pairs of projective surfaces

### 31 Mar 2026 v3

@@ 2 The Bourbaki degree of pairs of projective surfaces
In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials. We start with definitions and first results (Section 2.1), followed by a reduction to the case of a projective plane curve (Section 2.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (Section 2.3), in particular introducing the class of nearly free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
@@ 2.1 Framework and first results
and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$. We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [undefj].
The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
@@
Follows from direct application of the Hirzebruch-Riemann-Roch theorem. ∎
We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$. From the previous formula, $m(\sigma)$ is non-negative, and it is zero if and only if the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ is constant, that is, if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
As in the introduction, the Jacobian scheme of $\nabla\sigma$ is defined as
@@
the zero-locus of the $2\times 2$ minors of $\nabla\sigma$ or, in other words, the $0$-th Fitting scheme of $\nabla\sigma$. From the general theory of Fitting ideals (see, for example, [undefh, Chapter 20]), the annihilator ideal of $\mathcal{Q}_{\sigma}$ contains the Fitting ideal, and they have the same support, namely
$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
as closed subsets, and therefore the irreducible components of the schematic support $\operatorname{supp}(\mathcal{Q}_{\sigma})$ coincide with the irreducible components of $\Xi_{\sigma}$. In general, however, their schematic structure may be different, as we explore in the following example. Denote by $(\Xi_{\sigma})_{1}$ the one-dimensional part of the scheme $\Xi_{\sigma}$.
@@
$\begin{pmatrix}0&-2x_{1}+2x_{3}&0&2x_{1}\\ -3x_{1}x_{3}&3x_{1}^{2}-3x_{0}x_{3}&3x_{3}^{2}&-3x_{0}x_{1}+6x_{2}x_{3}\\ \end{pmatrix}$
and the matrix below
@@
gives two linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
In this case, the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$ and the 0-th Fitting ideal are different, given by:
$\begin{cases}\operatorname{Ann}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3},x_{0}x_{1}^{2}-x_{1}^{3})=(x_{1},x_{3})^{2}\cap(x_{3},x_{0}-x_{1})\\ \operatorname{Fitt}_{0}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3}^{2},x_{1}^{2}x_{3},x_{0}x_{1}^{2}-x_{1}^{3}-2x_{1}x_{2}x_{3}+2x_{2}x_{3}^{2}),\end{cases}$
so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively. From the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$, since ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ we conclude $m(\sigma)=5$. The right approach to computing $m(\sigma)$ in general is using the associativity formula, as we will in the next examples.
Any saturated syzygy of the Jacobian matrix $\nu$, of degree $e\in\mathbb{Z}$, induces a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described as the Serre correspondence (see [undefn, Theorem $4.1$]) between the sheaf ${\mathcal{T}}_{\sigma}$ and the scheme $B$. Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$. Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $e\geq 0$. The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $e$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
@@
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
For any saturated syzygy $\nu\in H^{0}({\mathcal{T}}(e))$ of degree $e\geq 0$, let $B_{\nu}\subset\mathbb{P}^{3}$ be the associated closed subscheme of pure dimension one. Then, we have the following equation:
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+m_{0}-m(\sigma),$
@@
$m(\sigma)$ $=\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})+\frac{d_{f}^{2}+d_{g}^{2}}{2}=\frac{c_{1}^{2}({\mathcal{T}}_{\sigma})-2c_{2}({\mathcal{T}}_{\sigma})}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
@@
On the other hand, since $c_{2}({\mathcal{T}}_{\sigma}(e))=\deg(B_{\nu})$ and ${\mathcal{T}}_{\sigma}$ is reflexive of rank $2$, we have
$\deg(B_{\nu})=c_{2}({\mathcal{T}}_{\sigma}(e))=c_{2}({\mathcal{T}}_{\sigma})+e\cdot c_{1}({\mathcal{T}}_{\sigma})+e^{2}.$
@@
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
We define the Bourbaki degree of a sequence $\sigma$, inspired by the construction in [undefr, Definition 2.4].
@@
Let $\sigma$ be a normal sequence and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. The Bourbaki degree of a normal sequence $\sigma$ is defined by:
@@
for some non-trivial syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. We note that every non-trivial syzygy of minimal degree is saturated.
@@
splits.
Remark 6.
From [undefm, Proposition 4.1], we have the formula
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
relating the third Chern class of ${\mathcal{T}}_{\sigma}$ and the discrete invariants of $B$, when $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $B$ is the zero locus of a non-zero section in $H^{0}({\mathcal{T}}_{\sigma}(e))$. Moreover, dualizing the sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B}(e-d)\rightarrow 0$
we conclude that the singular set of the sheaf ${\mathcal{T}}_{\sigma}$ is contained in $B$.
@@
    We have $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
@@
are a syzygies of degree $d_{f}+d_{g}$. Since there is at least one nonzero $2\times 2$ minor, from the hypothesis of algebraically independent, at least one of the syzygies $\nu_{0},\ldots,\nu_{3}$ is nonzero.
@@
For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [undefj, Lemma 2.7]. Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [undefj, Lemma 2.8]. The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
To show $(1)\Rightarrow(2)$, let $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. From the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=e(e-d)$. Since $\operatorname{Bour}(\sigma)\geq 0$ and from part (a) we have $e\leq d$, it follows that $e=0$ or $e=d$. Both cases mean $\operatorname{Bour}(\sigma)=0$ and the sequence is free. Since $c_{1}({\mathcal{T}}_{\sigma})=-d$ is additive, it follows that $e=0$, otherwise we would have $c_{1}({\mathcal{T}}_{\sigma})<-d$. ∎
@@
The bound obtained in Theorem A, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory. We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [undefj, Section 9]). Let $C$ be the one-dimensional part of the singular scheme of this foliation. From the formulas of discrete invariants in [undefa, Theorem 3.1] given for generic foliations,
@@
and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [undefv, Corollary 4.8].
From simple observations about the formula $\operatorname{Bour}(\sigma)$, we are able to obtain the following inequalities related to $\mu$-stability and freeness of the logarithmic sheaves.
@@
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$. Denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $d=d_{f}+d_{g}$. Then:
@@
      A minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
      $0\to\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{1})\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{2})\xrightarrow{\varphi}\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}(-d)\xrightarrow{\psi}{\mathcal{T}}_{\sigma}\to 0,$
      a Buchsbaum–Rim resolution, where
      $\psi=\begin{pmatrix}0&\Delta_{12}&\Delta_{13}&\Delta_{14}\\ -\Delta_{12}&0&\Delta_{23}&\Delta_{24}\\ -\Delta_{13}&-\Delta_{23}&0&\Delta_{34}\\ -\Delta_{14}&-\Delta_{24}&-\Delta_{34}&0\end{pmatrix},\quad\phi=\begin{pmatrix}\partial_{0}f&-\partial_{0}g\\[2.0pt] -\partial_{1}f&\partial_{1}g\\[2.0pt] \partial_{2}f&-\partial_{2}g\\[2.0pt] -\partial_{3}f&\partial_{3}g\end{pmatrix}$,
      and each $\Delta_{ij}$ is the $(i,j)$-minor of the matrix $\nabla\sigma$. In particular, $e=d$.
@@
The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)\geq 0$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$, from the inequality $e\leq d$. This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility. The converse also follows simply from the formula, assuming that $m(\sigma)=0$ and $e=d$.
To show the equivalences above, we start by noting that $m(\sigma)=0$ implies that the grade of the ideal of $(2\times 2)$-minors is three, and therefore the Buchsbaum–Rim complex is a minimal free resolution for $\nabla\sigma$. From this fact, we obtain that $e=d$. For the converse, since $e=d$ and the Buchsbaum–Rim complex is a minimal resolution, we obtain that the grade of the ideal of minors of $\nabla\sigma$ is at least three, hence $\dim(\mathcal{Q}_{\sigma})=0$ and thus $m(\sigma)=0$.
By construction, $m(\sigma)\geq 0$, and therefore for a given value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, the Bourbaki degree of $\sigma$ can be at most
@@
and this is a function $H=H(e)$ which attains its minimum at $e=d/2$. The function $H$ is decreasing on $e\in\{1,\ldots,d/2\}$, and if $e>d/2$, ${\mathcal{T}}_{\sigma}$ is $\mu$-stable, since $\mu({\mathcal{T}}_{\sigma})=-d/2$. Thus, the maximum value $\operatorname{Bour}(\sigma)$ in the range $e\in\{1,\ldots,d/2\}$ for any possible value of $m(\sigma)$ is $H(1)$, which is the expression above on the right-hand side. If $\operatorname{Bour}(\sigma)$ is higher than this, then $e>d/2$, and thus we obtain $(b)$.
With the same strategy as $(b)$, since the function $H(e)$ attains its minimum at $e=d/2$, for $m(\sigma)$ satisfying the inequality of the claim we obtain that
$\operatorname{Bour}(\sigma)=H(e)\geq H(d/2)=\frac{d^{2}}{4}-\frac{d^{2}}{2}+m_{0}-m(\sigma)>0,$
and therefore $\operatorname{Bour}(\sigma)\neq 0$, independently of the value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. ∎

### 27 Apr 2026 v4

@@ 2 The Bourbaki degree of pairs of projective surfaces
In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials. We start with definitions and first results (Section 2.1), followed by a reduction to the case of a projective plane curve (Section 2.2) and finish with results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (Section 2.3).
@@ 2.1 Framework and first results
and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$. A sequence $\sigma$ will be called free if ${\mathcal{T}}_{\sigma}$ splits as a direct sum of line bundles. We assume throughout that the sequences $\sigma$ satisfy the condition $\operatorname{{codim}}(\mathcal{Q}_{\sigma})\geq 2$, and in this case we call $\sigma$ a normal sequence. With this assumption, we obtain $\operatorname{{rk}}(\mathcal{Q}_{\sigma})=0$ and $c_{1}(\mathcal{Q}_{\sigma})=0$. The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
@@
This is a direct application of the Hirzebruch-Riemann-Roch Theorem. ∎
We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$. Since $m(\sigma)$ is the leading coefficient of a Hilbert polynomial, $m(\sigma)\geq 0$, and it is zero if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
The Jacobian scheme of $\nabla\sigma$ is defined as
@@
the zero-locus of the $2\times 2$ minors of $\nabla\sigma$ or, in other words, the $0$-th Fitting scheme of $\nabla\sigma$. From the general theory of Fitting ideals, the annihilator ideal $\operatorname{\operatorname{Ann}}\mathcal{Q}_{\sigma}$ of $\mathcal{Q}_{\sigma}$ contains the Fitting ideal $\operatorname{Fitt}_{0}(\nabla\sigma)$, and they have the same reduced support, namely
$V(\sqrt{\operatorname{\operatorname{Ann}}(\mathcal{Q}_{\sigma}})=\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=V(\operatorname{Fitt}_{0}(\nabla\sigma))_{\text{red}}=V(\sqrt{\operatorname{Fitt}_{0}(\nabla\sigma)})$
as closed subsets of $\mathbb{P}^{3}$. In particular, the irreducible components of the schematic support $\operatorname{supp}(\mathcal{Q}_{\sigma})$ coincide with the irreducible components of $\Xi_{\sigma}$. The respective schematic structures may be different, as we explore in the next example. Denote by $(\Xi_{\sigma})_{1}$ the one-dimensional part of the scheme $\Xi_{\sigma}$.
@@
$\begin{pmatrix}0&-2x_{1}+2x_{3}&0&2x_{1}\\ -3x_{1}x_{3}&3x_{1}^{2}-3x_{0}x_{3}&3x_{3}^{2}&-3x_{0}x_{1}+6x_{2}x_{3}\\ \end{pmatrix}$.
The annihilator ideal of $\operatorname{coker}(\nabla\sigma)$ and the 0-th Fitting ideal are different, namely given by:
$\begin{cases}\operatorname{\operatorname{Ann}}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3},x_{0}x_{1}^{2}-x_{1}^{3})=(x_{1},x_{3})^{2}\cap(x_{3},x_{0}-x_{1})\\ \operatorname{Fitt}_{0}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3}^{2},x_{1}^{2}x_{3},x_{0}x_{1}^{2}-x_{1}^{3}-2x_{1}x_{2}x_{3}+2x_{2}x_{3}^{2}),\end{cases}$
so both schemes $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ are non-reduced, with degrees 44 and $6$, respectively. The matrix below
@@
gives two linearly independent syzygies for the matrix $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ and $\sigma$ is free. From this and the additivity of the Hilbert polynomial on short exact sequences, one may compute $m(\sigma)=5$.
The following proposition gives a sufficient condition to forbid the anomaly above to occur.
@@
Let $\sigma=(f,g)$ be a normal sequence. If $\operatorname{{codim}}(\operatorname{Fitt}_{1}(\nabla\sigma))\geq 3$, then
$m(\sigma)=\deg(\Xi_{\sigma})=\deg(\operatorname{supp}\mathcal{Q}_{\sigma}).$
Proof.
Let $Q_{\sigma}$ denote the graded $R$-module associated to $\mathcal{Q}_{\sigma}$. All three invariants above may be described using the associativity formula, as follows:
$\deg(\Xi)$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(\operatorname{Fitt}_{0}(\nabla\sigma)),\;\dim(R/\mathfrak{p})=\dim(\operatorname{Fitt}_{0}(\nabla\sigma)_{\mathfrak{p}})}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}\bigl((R/\operatorname{Fitt}_{0}(\nabla\sigma))_{\mathfrak{p}}\bigr)\cdot\deg(R/\mathfrak{p})$
$\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(\operatorname{Fitt}_{0}(\nabla\sigma)),\;\dim(R/\mathfrak{p})=\dim((Q_{\sigma})_{\mathfrak{p}})}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}\bigl((R/\operatorname{\operatorname{Ann}}Q_{\sigma})_{\mathfrak{p}}\bigr)\cdot\deg(R/\mathfrak{p})$
$m(\sigma)$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(\operatorname{Fitt}_{0}(\nabla\sigma)),\;\dim(R/\mathfrak{p})=\dim((Q_{\sigma})_{\mathfrak{p}})}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}\bigl((Q_{\sigma})_{\mathfrak{p}}\bigr)\cdot\deg(R/\mathfrak{p}).$
The sets of primes involved in each sum are the same, so we only need to check that the lengths coincide over minimal prime ideals $\mathfrak{p}$ of height two. Since $\operatorname{{codim}}(\operatorname{Fitt}_{1}(\nabla\sigma))\geq 3$, the rank of the localization $(\nabla\sigma)_{\mathfrak{p}}$ is one, and we may change coordinates locally to write
$A_{\mathfrak{p}}\doteq(\nabla\sigma)_{\mathfrak{p}}=\begin{pmatrix}1&0&0&0\\ 0&a&b&c\end{pmatrix}$.
In these coordinates, we obtain $\operatorname{Fitt}_{0}(A_{\mathfrak{p}})=(a,b,c)$ from the $2\times 2$-minors, and moreover $(Q_{\sigma})_{\mathfrak{p}}\simeq R_{\mathfrak{p}}/(a,b,c)$. We claim that $\operatorname{\operatorname{Ann}}(Q_{\sigma})_{\mathfrak{p}}=(a,b,c)$. Since $(a,b,c)=\operatorname{Fitt}_{0}(\nabla\sigma)\subset\operatorname{\operatorname{Ann}}(Q_{\sigma})$, we show the other inclusion. Let $r\in R_{\mathfrak{p}}$ such that $r(Q_{\sigma})_{\mathfrak{p}}=0$. Writing $(Q_{\sigma})_{\mathfrak{p}}=\operatorname{coker}(A_{\mathfrak{p}})$, we have
$r\cdot\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}=\begin{pmatrix}rv_{1}\\ rv_{2}\end{pmatrix}\in\text{im }A_{\mathfrak{p}}$
for $(v_{1},v_{2})\in R_{\mathfrak{p}}^{2}$. From the second equation, one obtains $r\cdot v_{2}=au_{1}+bu_{2}+cu_{3}$, for some $u_{i}\in A_{p}$. Taking $v_{2}=1$ yields $r\in(a,b,c)$. This shows that
$\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}}/\operatorname{Fitt}_{0}(\nabla\sigma)_{\mathfrak{p}})=\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}}/\operatorname{\operatorname{Ann}}(Q_{\sigma})_{\mathfrak{p}})=\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}((Q_{\sigma})_{\mathfrak{p}}),$
and since $(Q_{\sigma})_{\mathfrak{p}}\simeq R_{\mathfrak{p}}/(a,b,c)$, the claim follows. ∎
Given a sequence $\sigma=(f,g)$ and a saturated syzygy of the Jacobian matrix $\nu$ of degree $l\in\mathbb{Z}$, we obtain a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated with $\nu$. Using the Hilbert polynomial, one can write formulas (see [14, Proposition $4.1$]):
$\deg(B_{\nu})$ $=c_{2}({\mathcal{T}}_{\sigma}(l))$
$c_{3}({\mathcal{T}}_{\sigma})$ $=2p_{a}(B)-2+\deg(B)(4+d-2l).$
Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $l\geq 0$. The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $l$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
Proposition 4.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$. For a saturated syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(l))$ of degree $l\geq 0$, let $B_{\nu}\subset\mathbb{P}^{3}$ be the associated curve. Then, we have the following equation:
$\deg(B_{\nu})=l^{2}-l(d_{f}+d_{g})+m_{0}-m(\sigma),$
@@
$m(\sigma)$ $=\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
$=\frac{c_{1}^{2}({\mathcal{T}}_{\sigma})-2c_{2}({\mathcal{T}}_{\sigma})}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
@@
On the other hand, since $c_{2}({\mathcal{T}}_{\sigma}(l))=\deg(B_{\nu})$ and ${\mathcal{T}}_{\sigma}$ is reflexive of rank $2$, we have
$\deg(B_{\nu})=c_{2}({\mathcal{T}}_{\sigma}(l))=c_{2}({\mathcal{T}}_{\sigma})+l\cdot c_{1}({\mathcal{T}}_{\sigma})+l^{2}.$
@@
$\deg(B_{\nu})=l^{2}-l(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
We define the Bourbaki degree of a sequence $\sigma$, inspired by the construction in [17, Definition 2.4].
@@
Let $\sigma=(f,g)$ be a normal sequence, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $d\doteq d_{f}+d_{g}$. The Bourbaki degree of $\sigma$ is defined by:
@@
for some non-trivial syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. We note that every non-trivial syzygy of minimal degree is saturated, and $\operatorname{Bour}(\sigma)$ is independent of the choice of $\nu$.
@@
splits, as the extension group $\operatorname{{Ext}}^{1}(\mathcal{O}_{\mathbb{P}^{3}}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(d-2e))$ vanishes.
Whenever, up to a change of coordinates, the elements of a sequence $\sigma=(f,g)$ do not depend on some variable in $R$, $\sigma$ is called compressible, see [11]. The following lemma is a consequence of their results in our context:
Lemma 7.
[[11, Lemmas 2.7 and 2.8]] Let $\sigma=(f,g)$ be a normal sequence in $\mathbb{P}^{3}$. Then:
  • (a)
    A sequence $\sigma$ is compressible if and only if $h^{0}({\mathcal{T}}_{\sigma})\neq 0$;
  • (b)
    If $\sigma$ is compressible, then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
@@
    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
@@
are a syzygies of degree $d_{f}+d_{g}$. Since $\sigma$ is algebraically independent, there is at least one nonzero $2\times 2$ minor, and at least one of the syzygies $\nu_{0},\ldots,\nu_{3}$ is nonzero.
@@
For $(c)$, the equivalence $(2)\iff(3)$ is the content of 7, $(a)$. Moreover, $(3)\Rightarrow(4)$ is 7, $(b)$. The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
To show $(1)\Rightarrow(2)$, let $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. From the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=e(e-d)$. Since $\operatorname{Bour}(\sigma)\geq 0$ and, from part (a), $e(e-d)\leq 0$, it follows that either $e=0$ or $e=d$. Both cases mean $\operatorname{Bour}(\sigma)=0$ and the sequence is free, say ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-a)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-b)$, with $0\leq a\leq b$. If $e=d$, then $a=d>0$, but with $c_{1}({\mathcal{T}}_{\sigma})=-(a+b)=-d$ one gets a contradiction with $a\leq b$, and thus $e=0$. ∎
@@
The bound obtained in Theorem A, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory. We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension-one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [11, Section 9]). Let $C$ be the one-dimensional part of the singular scheme of this foliation. From the formulas of discrete invariants in [3, Theorem 3.1]:
@@
and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [21, Corollary 4.8].
The following inequalities relate the discrete invariants with $\mu$-stability and freeness of the logarithmic sheaves.
@@
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$. Denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $d=d_{f}+d_{g}$. Then:
@@
      ${\mathcal{T}}_{\sigma}$ admits a minimal resolution by the Buchsbaum–Rim complex:
      $0\to\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{f})\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{g})\xrightarrow{\varphi}\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}(-d)\xrightarrow{\psi}{\mathcal{T}}_{\sigma}\to 0,$
      where
      $\psi=\begin{pmatrix}0&\Delta_{12}&\Delta_{13}&\Delta_{14}\\ -\Delta_{12}&0&\Delta_{23}&\Delta_{24}\\ -\Delta_{13}&-\Delta_{23}&0&\Delta_{34}\\ -\Delta_{14}&-\Delta_{24}&-\Delta_{34}&0\end{pmatrix},\quad\varphi=\begin{pmatrix}\partial_{0}f&-\partial_{0}g\\[2.0pt] -\partial_{1}f&\partial_{1}g\\[2.0pt] \partial_{2}f&-\partial_{2}g\\[2.0pt] -\partial_{3}f&\partial_{3}g\end{pmatrix}$,
      and each $\Delta_{ij}$ is the $(i,j)$-minor of the matrix $\nabla\sigma$. In particular, $e=d$ and the logarithmic tangent sheaf is identified with the cokernel of the transpose matrix ${\mathcal{T}}_{\sigma}(d)\simeq\operatorname{coker}(\nabla\sigma)^{T}$.
@@
The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)=e(e-d)+m_{0}-m(\sigma)$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$. This also shows equality occurs if and only if both e(ed)e(e-d) and $m(\sigma)$ are zero, therefore $e=d$ or $e=0$. But $e=0$ gives $\operatorname{Bour}(\sigma)=0$, from compressibility, and $m_{0}\neq 0$ implies $e=d$. To show the equivalences above, note that $m(\sigma)=0$ if and only if the grade of the ideal of $(2\times 2)$-minors is three, which occurs if and only if the Buchsbaum–Rim complex is a minimal free resolution for $\nabla\sigma$, when in particular $e=d$.
To show $(b)$, note that
@@
and this is a function $H=H(e)$ which attains its minimum at $e=d/2$. The function $H$ is decreasing on $e\in\{1,\ldots,d/2\}$, so when $\operatorname{Bour}(\sigma)>H(1)=(d_{f}-1)(d_{f}+d_{g})+d_{g}^{2}+1$, then $e>d/2$, and thus ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
To prove $(c)$, note that $H$ attains its minimum at $e=d/2$, so we obtain that
$\operatorname{Bour}(\sigma)=H(e)-m(\sigma)\geq H(d/2)-m(\sigma)=\frac{d^{2}}{4}-\frac{d^{2}}{2}+m_{0}-m(\sigma)>0,$
and therefore $\operatorname{Bour}(\sigma)\neq 0$, independently of the value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and $\sigma$ is not free. ∎